# Improving Automated Feedback Systems for Tutor Training in Low-Resource Scenarios through Data Augmentation

### v1

@@ Abstract
Tutoring is an effective instructional method for enhancing student learning, yet its success relies on the skill and experience of the tutors. This reliance presents challenges for the widespread implementation of tutoring, particularly in training novice tutors. To support tutor training programs, real-time automated feedback systems are essential for efficiently training large numbers of tutors. Lin et al.’s previous study employed Generative Pre-Trained Transformers (GPT) for sequence labeling to identify desirable and undesirable praise components in a tutor training dataset, providing explanatory feedback. However, this approach requires a significant amount of labeled data for fine-tuning, which is both labor-intensive and dependent on expert input. To address the challenges associated with extensive data labeling, the current study explores the use of prompting more advanced GPT models like GPT-4o to generate synthetic datasets for augmenting labeled response data, followed by fine-tuning a GPT-3.5 model. Our results demonstrate that: (1) fine-tuning with augmented dataset with size 520 generated by GPT-4o significantly improves GPT-3.5’s performance in identifying praise components, with $F_{2}$ increasing by 17.8% for effort-based praise and 20.8% for outcome-based praise compared to the same model fine-tuned without augmentation. These improvements are also reflected in traditional metrics like M-IoUand IoU score. (2) Our data augmentation approach generalizes effectively to identify other types of praise (i.e., person-based praise), with $F_{2}$ Score increasing by 20.4% for person-based praise, alongside a 19.9% increase in M-IoU and a 21.6% increase in IoU with augmented dataset size 520, compared to the same model fine-tuned without augmentation. These findings suggest that for data-intensive tasks, synthetic data generated through GPT model prompting can substantially enhance fine-tuned model performance in low-resource scenarios.
Index Terms:
Tutor training, large language models, data augmentation, sequence labeling, automated feedback system.
@@ I Introduction
Human tutoring is an effective instructional method known for its efficacy in enhancing student learning. However, several logistical challenges hinder its widespread implementation, including the recruitment, training, and retention of skilled tutors [1]. Training tutors is resource-intensive and often requires hands-on mentorship from experienced tutors [2]. A key aspect of effective tutor training involves equipping novice tutors with proficient tutoring strategies [3, 1]. For example, rather than merely correcting an incorrect answer, skilled tutors engage students to uncover underlying misconceptions, enabling more effective support. Traditionally, these insights are imparted through direct training from experienced tutors, but this approach is limited by the shortage of expert tutors and the high cost of personalized mentorship. [3, 4, 5, 6].
In response to the increasing demand for scalable hands-on support in tutor training, researchers are increasingly leveraging automated feedback systems [7, 8, 6, 9, 10, 11, 12]. While many implementations utilize AI algorithms to generate automated feedback [13], their specific application to tutor training remains under-explored. This gap is significant because AI has the potential to enhance scalability and effectiveness in tutor training, addressing the current limitations of hands-on tutor training systems that are resource-intensive and difficult to scale. Traditional models like BERT [14] face challenges due to limited labeled training datasets and limited generalization ability, which constrain their ability to provide precise feedback [3, 15]. Recent advances in Large Language Models (LLMs) offer a promising solution (e.g.,[6]) to these challenges like limited labeled training datasets and bad generalization ability due to their strong generalization capabilities and vast knowledge base [16, 17]. These LLMs can dynamically adapt to domain-specific scenarios like generating educational feedback, making them well-suited for developing real-time feedback systems for tutor training [6, 16].
Although fine-tuning these LLM models has shown high accuracy in generating educational feedback, this approach requires a substantial amount of labeled response data, which is both labor-intensive and often requires specialized expertise [6]. This challenge is especially noticeable in situations with limited data (low-resource scenarios), which makes the fine-tuning process more difficult. A promising method to address the challenge of low-resource data is text data augmentation using natural language processing techniques, which can generate synthetic data samples and expand the training dataset.
@@ II-C Sequence Labeling for Feedback Generation
Recent advancements in NLP have underscored the potential of LLMs, such as GPT-3.5 and GPT-4 [38, 39], in various educational contexts through techniques like prompting and fine-tuning [40]. These models have demonstrated significant promise in enhancing the identification and categorization of key text segments [41, 38, 39], which is critical for developing automated systems that provide targeted, explanatory feedback. Our previous work [6] investigated the use of prompting and fine-tuning GPT models to identify both desired and undesired components in tutor open-ended responses involving praise. This study [6] evaluated the models’ capabilities in developing an automated system for providing explanatory feedback to tutors, and found that fine-tuning significantly outperformed prompting. Fine-tuning GPT models has shown considerable potential in educational applications [40]. Fine-tuning involves adjusting the model’s neural network to better fit specific domains, thereby enhancing its performance in those contexts [36]. Building on our prior work, we aim to further enhance model performance in low-resource scenarios by leveraging data augmentation techniques, thereby extending our previous findings and methods.
@@ III-A Dataset
Our study received ethical approval from the Institutional Review Board (IRB) at Carnegie Mellon University. For RQ1, we used the same dataset as our previous study [6] to maintain consistency and facilitate comparison. This dataset includes responses from 65 volunteer tutors who participated in the Giving Effective Praise lesson. The demographic breakdown of these tutors was: 52% identified as White, 18% as Asian, 52% were male, and over half were aged 50 years or older. The purpose of the Giving Effective Praise lesson is to instruct tutors on the effective use of praise to boost student motivation. We collected a total of 129 responses from tutors who completed the lesson, categorizing each response by the type of praise used (i.e., effort-based praise and outcome-based praise). In our previous work, we divided the entire dataset of 129 labeled responses into a training set containing 65 labeled responses and a test set containing 64 labeled responses. For the current study, we focus on exploring low-resource scenarios by using 10% of the total dataset, namely 13 labeled responses, as the new training set. This extreme reduction is intended to simulate real-world situations where the availability of labeled data is severely limited. We then apply data augmentation techniques to this reduced subset. The choice of 10% was made to ensure that there is sufficient initial information available to guide the augmentation process. Using a smaller subset would severely limit the semantic richness of the initial data, making it difficult for the augmentation methods to generate diverse and meaningful examples. We aim to strike a balance between simulating a low-resource scenario and providing enough foundational data to enable the augmentation techniques to produce useful, varied, and contextually appropriate examples.
@@ III-C1 Fine-tuning GPTs
We utilized GPT-3.5 (gpt-3.5-turbo-1106) to evaluate the effectiveness of data augmentation. This approach involved fine-tuning the GPT-3.5 model to recognize and understand patterns related to identifying praise components in tutor responses. To prepare the data for fine-tuning, we converted the tutor responses and their corresponding tags into JSON format. This structured format aligns with the input style required by the GPT model. The details of our fine-tuning scheme can be found in Appendix A.
@@ III-C2 Data Augmentation with GPT-4o
To address the issue of low-resource in fine-tuning GPT-3.5 for tutor response scenarios, we leveraged more advanced model GPT-4o (gpt-4o) to perform data augmentation and increase the number of labeled responses. Our initial attempt to use GPT-4o directly for data augmentation revealed that the model tended to generate easy and typical responses and labels, which limited the fine-tuned model’s ability to learn from difficult cases. This issue arises because generalized LLMs often produce responses that lack the diversity and complexity needed for nuanced educational feedback tasks. As a result, these typical outputs do not adequately challenge the model to improve its performance on more complex or less common feedback scenarios. To overcome these limitations, we developed a structured data augmentation approach using GPT-4o. Our approach focuses on generating synthetic datasets that retain authentic sentence structures and expressions, which ensures that the augmented data is both semantically coherent and contextually relevant, providing a richer and more varied training set. By doing this, we aim to introduce variability while preserving the grammaticality and coherence of the text. Such targeted data augmentation leads to improved model performance by providing the student LLMs (i.e., GPT-3.5) with more relevant and contextually rich examples, ultimately resulting in more accurate and reliable performance.
@@
The prompt we used is shown in Table I. The TEXT represents different components of the sentence to be augmented, including effort-based praise, outcome-based praise, person-based praise, and other non-praise components, as illustrated in Figure 2. To ensure the quality of the generated synonyms, we set the temperature parameter to 0. The temperature setting controls the variability of the model’s output, and a value of 0 promotes more deterministic and high-quality responses [54]. During our prompt engineering process, we experimented with generating 5, 10, 15, 20, 25, and 30 synonyms. We found that generating 15 synonyms was optimal for ChatGPT-4o to produce high-quality synonyms without resorting to low-frequency words. For example, generating more than 15 synonyms for “You did a good job” sometimes resulted in rare expressions like “You executed a commendable operation”, which, while technically correct, is not commonly used in everyday educational settings. We tested five different prompts and observed consistent results across them. The quality of synthetic data is crucial, as demonstrated by our experiments with traditional data augmentation methods, which sometimes produced low-quality outputs (see Section 4.1.2 for results). In praise type labeling, even a single word difference can change the classification from effort-based to outcome-based praise. For instance, “You did a great job” (outcome-based) versus “You did a great job persevering” (effort-based) highlights how nuanced differences affect categorization. Therefore, high-quality data is essential for accurate model training.
TABLE I: Prompt for Data Augmentation
@@ III-D Metrics
In sequence labeling tasks, traditional metrics like the $F_{1}$ score are commonly used to assess model performance [55]. However, the inclusion of additional entities did not hinder the trainees’ ability to understand the correctness of the responses. For example, Table II illustrates that while expert annotations capture outcome-based praise accurately, model-generated annotations may include extra words (FP) that are still useful, unlike incomplete annotations (FN) that miss the praise’s intent. This suggests a need for a metric that more flexibly accommodates additional praise tokens. Therefore, we propose modifying the Intersection over Union (IoU) concept, commonly used in computer vision, to better suit our evaluation needs.
TABLE II: Original and predicted praise with outcome-based praise in red and effort-based praise in blue.
@@
The IoU metric (Equation 1), used in object detection and segmentation tasks, measures the overlap between predicted and actual annotations [56, 57]. In sequence labeling, the ‘Area of Overlap’ (TP) includes tokens correctly identified as praise, while the ‘Area of Union’ includes all tokens labeled as praise by the model (TP and FP) and all actual praise tokens in the groundtruth (TP and FN). The IoU metric inspired us in handling the challenge of additional detected words. To address additional detected words, we propose a metric that more flexibly accommodates additional praise tokens: Modified IoU (M-IoU) metric (Equation 2) that includes a weight coefficient, $\alpha$, reducing the impact of FPs and adding flexibility. The $\alpha$ value, set at 0.2 based on expert observations, allows adjustment of tolerance for additional words, which equates to a penalty level where five extra words are considered equivalent to one missing word. This reflects our approach’s emphasis on being more tolerant of extra words while being less forgiving of missing words. For instance, if an output contains five unnecessary words, it incurs a penalty similar to having one crucial word missing, aligning with our dataset’s annotation guidelines where annotators are instructed to mark the minimal words necessary to represent praise. As shown in Table II, including more extra words does not significantly affect the praise itself, whereas missing even one word can render it invalid. In cases where no praise is present and the model agrees (i.e., TP + FP + FN equals 0), we assign a score of 1 to reflect perfect agreement, showcasing the M-IoU’s adaptability and effectiveness in practical applications. In our previous work [6], we validated the effectiveness of the M-IoU score for evaluating the quality of highlighted components from GPT models through correlation analysis. We also acknowledge that the $F_{2}$ score (Equation 3) can achieve a similar effect by assigning a weight of 0.2 to FP and 0.8 to FN. However, unlike M-IoU, the $F_{2}$ score does not provide the flexibility to adjust the balance between penalties for FP and FN tokens. For instance, with the $F_{2}$ score, approximately four extra words (FP) incur the same penalty as one missing word (FN), and this ratio is fixed. In contrast, M-IoU allows us to fine-tune this balance by adjusting the $\alpha$ coefficient. However, since the $F_{2}$ score is close to our M-IoU setup here ($\alpha=0.2$), we have included it in our analysis as an additional metric to provide further insights.
IoU=Area of OverlapArea of $\text{IoU}=\frac{\text{Area of Overlap}}{\text{Area of Union}}=\frac{TP}{TP+FP+FN}$ (1)
[…]
@@ IV-A1 Experimental Setup
We begin by randomly extracting 13 labeled responses from the training set (consists of 65 labeled responses) to create a low-resource training set. This low-resource training set is then augmented using ChatGPT-4o under various data augmentation multipliers. These multipliers determine how many times the original number of responses is increased. For example, if the multiplier is 3 and we start with 5 responses, the augmented dataset will contain 15 responses. We evaluate the fine-tuned model’s performance on the test set consisting of 64 labeled responses. This experiment is repeated five times using five different random seeds when extracting the 13 labeled responses, and the average M-IoU performance is calculated. It should be noted that the five random seeds are the same as those used in our previous work [6].

### v2

@@ Abstract
Tutoring is an effective instructional method for enhancing student learning, yet its success relies on the skill and experience of the tutors. This reliance presents challenges for the widespread implementation of tutoring, particularly in training novice tutors. To support tutor training programs, real-time automated feedback systems are essential for efficiently training large numbers of tutors. Lin et al.’s previous study employed Generative Pre-Trained Transformers (GPT) for sequence labeling to identify desirable and undesirable praise components in a tutor training dataset, providing explanatory feedback. However, this approach requires a significant amount of labeled data for fine-tuning, which is both labor-intensive and dependent on expert input. To address the challenges associated with extensive data labeling, the current study explores the use of prompting more advanced GPT models like GPT-4o to generate synthetic datasets for augmenting labeled response data, followed by fine-tuning a GPT-3.5 model. Our results demonstrate that: (1) fine-tuning with augmented dataset with size 520 generated by GPT-4o significantly improves GPT-3.5’s performance in identifying praise components, with $F_{2}$ increasing by 20.5% for effort-based praise and 17.3% for outcome-based praise compared to the same model fine-tuned without augmentation. These improvements are also reflected in traditional metrics like M-IoUand IoU score. (2) Our data augmentation approach generalizes effectively to identify other types of praise (i.e., person-based praise), with $F_{2}$ Score increasing by 16.5% for person-based praise, alongside a 19.9% increase in M-IoU and a 21.6% increase in IoU with augmented dataset size 520, compared to the same model fine-tuned without augmentation. These findings suggest that for data-intensive tasks, synthetic data generated through GPT model prompting can substantially enhance fine-tuned model performance in low-resource scenarios. This approach offers a scalable and cost-effective solution to alleviate the reliance on extensive expert-labeled data, enabling broader adoption of automated tutor training systems and facilitating the rapid development of skilled tutors at scale.
@@ I Introduction
Human tutoring is an effective instructional method known for its efficacy in enhancing student learning. However, several logistical challenges hinder its widespread implementation, including the recruitment, training, and retention of skilled tutors [1]. Training tutors is resource-intensive and often requires hands-on mentorship. Equipping novice tutors with effective tutoring strategies is a key aspect of this training. For example, rather than merely correcting an incorrect answer, skilled tutors engage students to uncover underlying misconceptions, enabling more effective support. Traditionally, these insights are imparted through direct training from experienced tutors, but this approach is limited by the shortage of expert tutors and the high cost of personalized mentorship  [2, 3, 4, 5].
In response to the increasing demand for scalable hands-on support in tutor training, researchers are increasingly leveraging automated feedback systems [6, 7, 5, 8, 9, 10, 11]. Many AI algorithms generate automated feedback [12], but their specific use in tutor training remains under-explored. AI could improve scalability and effectiveness, addressing the limitations of resource-intensive hands-on training. Traditional models like BERT [13] face challenges due to limited labeled training datasets and limited generalization ability, which constrain their ability to provide precise feedback [2, 14]. Recent advances in Large Language Models (LLMs) offer a promising solution (e.g.,[5]) to challenges such as limited labeled training datasets and poor model robustness by leveraging their vast knowledge base and strong generalization capabilities [15, 16]. These LLMs can dynamically adapt to domain-specific scenarios like generating educational feedback, making them well-suited for developing real-time feedback systems for tutor training [5, 15].
Fine-tuning LLMs yields accurate educational feedback but demands large amounts of labeled data, which is labor-intensive and requires specialized expertise [5]. This challenge is especially noticeable in situations with limited data (low-resource scenarios), which makes the fine-tuning process more difficult. A promising method to address the challenge of low-resource data is text data augmentation using natural language processing techniques, which can generate synthetic data samples and expand the training dataset.
@@ II-C Sequence Labeling for Feedback Generation
Recent advancements in NLP have underscored the potential of LLMs, such as GPT-3.5 and GPT-4 [37, 38], in various educational contexts through techniques like prompting and fine-tuning [39]. These models have demonstrated significant promise in enhancing the identification and categorization of key text segments [40, 37, 38], which is critical for developing automated systems that provide targeted, explanatory feedback. Our previous work [5] investigated the use of prompting and fine-tuning GPT models to identify both desired and undesired components in tutor open-ended responses involving praise. This study [5] evaluated the models’ capabilities in developing an automated system for providing explanatory feedback to tutors, and found that fine-tuning significantly outperformed prompting. Fine-tuning GPT models has shown considerable potential in educational applications [39]. Fine-tuning involves adjusting the model’s neural network to better fit specific domains, thereby enhancing its performance in those contexts [35]. Building on our prior work, we aim to further enhance model performance in low-resource scenarios by leveraging data augmentation techniques, extending our previous findings and methods.
@@ III-A Dataset
Our study received ethical approval from the Institutional Review Board (IRB) at Carnegie Mellon University. For RQ1, we used the same dataset as our previous study [5] to maintain consistency and facilitate comparison. This dataset includes responses from 65 volunteer tutors who participated in the Giving Effective Praise lesson. The demographic breakdown of these tutors was: 52% identified as White, 18% as Asian, 52% were male, and over half were aged 50 years or older. The purpose of the Giving Effective Praise lesson is to instruct tutors on the effective use of praise to boost student motivation. We collected a total of 129 responses from tutors who completed the lesson, categorizing each response by the type of praise used (i.e., effort-based praise and outcome-based praise). In our previous work, we partitioned a dataset of 129 labeled responses into a training set of 64 responses (44 containing effort-based praise and 23 containing outcome-based praise) and a test set of 65 responses (47 containing effort-based praise and 33 containing outcome-based praise). In this study, we examine low-resource settings by utilizing only 10% of the entire dataset—namely, 13 labeled responses. Notably, all 13 responses in this reduced training set include both effort-based and outcome-based praise. The dataset details are also shown in Table I. This extreme reduction is intended to simulate real-world situations where the availability of labeled data is severely limited. We then apply data augmentation techniques to this reduced subset. The choice of 10% was made to ensure that there is sufficient initial information available to guide the augmentation process. Using a smaller subset would severely limit the semantic richness of the initial data, making it difficult for the augmentation methods to generate diverse and meaningful examples. We aim to strike a balance between simulating a low-resource scenario and providing enough foundational data to enable the augmentation techniques to produce useful, varied, and contextually appropriate examples.
TABLE I: Summary Statistics of Datasets for RQ 1
Research Question Dataset Split Total Responses Effort-based Outcome-based
RQ1 Training Set (13 subset) 13 13 13
Test Set 65 47 33
RQ1 full Training Set 64 44 23
Test Set 65 47 33
@@ III-C1 Fine-tuning GPTs
We utilized GPT-3.5 to evaluate the effectiveness of data augmentation. This approach involved fine-tuning the GPT-3.5 model to recognize and understand patterns related to identifying praise components in tutor responses. To prepare the data for fine-tuning, we converted the tutor responses and their corresponding tags into JSON format. This structured format aligns with the input style required by the GPT model. The details of our fine-tuning scheme can be found in Appendix A.
@@ III-C2 Data Augmentation with GPT-4o
To address the issue of low-resource in fine-tuning GPT-3.5 for tutor response scenarios, we leveraged more advanced model GPT-4o to perform data augmentation and increase the number of labeled responses. Our initial attempt to use GPT-4o directly for data augmentation revealed that the model tended to generate easy and typical responses and labels, which limited the fine-tuned model’s ability to learn from difficult cases. This issue arises because generalized LLMs often produce responses that lack the diversity and complexity needed for nuanced educational feedback tasks. As a result, these typical outputs do not adequately challenge the model to improve its performance on more complex or less common feedback scenarios. To overcome these limitations, we developed a structured data augmentation approach using GPT-4o. Our approach focuses on generating synthetic datasets that retain authentic sentence structures and expressions, which ensures that the augmented data is both semantically coherent and contextually relevant, providing a richer and more varied training set. By doing this, we aim to introduce variability while preserving the grammaticality and coherence of the text. Such targeted data augmentation leads to improved model performance by providing the student LLMs (i.e., GPT-3.5) with more relevant and contextually rich examples, ultimately resulting in more accurate and reliable performance.
@@
The prompt we used is shown in Table II. The TEXT represents different components of the sentence to be augmented, including effort-based praise, outcome-based praise, person-based praise, and other non-praise components, as illustrated in Figure 2. To ensure the quality of the generated synonyms, we set the temperature parameter to 0. The temperature setting controls the variability of the model’s output, and a value of 0 promotes more deterministic and high-quality responses [53]. During our prompt engineering process, we experimented with generating 5, 10, 15, 20, 25, and 30 synonyms. We found that generating 15 synonyms was optimal for ChatGPT-4o to produce high-quality synonyms without resorting to low-frequency words. For example, generating more than 15 synonyms for “You did a good job” sometimes resulted in rare expressions like “You executed a commendable operation”, which, while technically correct, is not commonly used in everyday educational settings. We tested five different prompts and observed consistent results across them. The quality of synthetic data is crucial, as demonstrated by our experiments with traditional data augmentation methods, which sometimes produced low-quality outputs (see Section 4.1.2 for results). In praise type labeling, even a single word difference can change the classification from effort-based to outcome-based praise. For instance, “You did a great job” (outcome-based) versus “You did a great job persevering” (effort-based) highlights how nuanced differences affect categorization. Therefore, high-quality data is essential for accurate model training.
TABLE II: Prompt for Data Augmentation
@@ III-D Metrics
In sequence labeling tasks, traditional metrics like the $F_{1}$ score are commonly used to assess model performance [54]. However, the inclusion of additional entities did not hinder the trainees’ ability to understand the correctness of the responses. For example, Table III illustrates that while expert annotations capture outcome-based praise accurately, model-generated annotations may include extra words (FP) that are still useful, unlike incomplete annotations (FN) that miss the praise’s intent. This suggests a need for a metric that more flexibly accommodates additional praise tokens. Therefore, we propose modifying the Intersection over Union (IoU) concept, commonly used in computer vision, to better suit our evaluation needs.
TABLE III: Original and predicted praise with outcome-based praise in red and effort-based praise in blue.
@@
The IoU metric (Equation 1), used in object detection and segmentation tasks, measures the overlap between predicted and actual annotations [55, 56]. In sequence labeling, the ‘Area of Overlap’ (TP) includes tokens correctly identified as praise, while the ‘Area of Union’ includes all tokens labeled as praise by the model (TP and FP) and all actual praise tokens in the groundtruth (TP and FN). The IoU metric inspired us in handling the challenge of additional detected words. To address additional detected words, we propose a metric that more flexibly accommodates additional praise tokens: Modified IoU (M-IoU) metric (Equation 2) that includes a weight coefficient, $\alpha$, reducing the impact of FPs and adding flexibility. The $\alpha$ value, set at 0.2 based on expert observations, allows adjustment of tolerance for additional words, which equates to a penalty level where five extra words are considered equivalent to one missing word. This reflects our approach’s emphasis on being more tolerant of extra words while being less forgiving of missing words. For instance, if an output contains five unnecessary words, it incurs a penalty similar to having one crucial word missing, aligning with our dataset’s annotation guidelines where annotators are instructed to mark the minimal words necessary to represent praise. As shown in Table III, including more extra words does not significantly affect the praise itself, whereas missing even one word can render it invalid. In cases where no praise is present and the model agrees (i.e., TP + FP + FN equals 0), we assign a score of 1 to reflect perfect agreement, showcasing the M-IoU’s adaptability and effectiveness in practical applications. In our previous work [5], we validated the effectiveness of the M-IoU score for evaluating the quality of highlighted components from GPT models through correlation analysis. We also acknowledge that the $F_{2}$ score (Equation 3) can achieve a similar effect by assigning a weight of 0.2 to FP and 0.8 to FN. However, unlike M-IoU, the $F_{2}$ score does not provide the flexibility to adjust the balance between penalties for FP and FN tokens. For instance, with the $F_{2}$ score, approximately four extra words (FP) incur the same penalty as one missing word (FN), and this ratio is fixed. In contrast, M-IoU allows us to fine-tune this balance by adjusting the $\alpha$ coefficient. However, since the $F_{2}$ score is close to our M-IoU setup here ($\alpha=0.2$), we have included it in our analysis as an additional metric to provide further insights.
$\text{IoU}=\frac{\text{Area of Overlap}}{\text{Area of Union}}=\frac{TP}{TP+FP+FN}$ (1)
$\text{M-IoU}=\frac{TP}{TP+\alpha\times FP+FN}$ (2)
$\begin{split}\text{$F_{2}$ score}&=\frac{(1+2^{2})\times TP}{(1+2^{2})\times TP+FP+2^{2}\times FN}\\ &=\frac{TP}{TP+0.2\times FP+0.8\times FN}\end{split}$ (3)
@@ IV-A1 Experimental Setup
We begin by randomly extracting 13 labeled responses from the training set (consists of 65 labeled responses) to create a low-resource training set. This low-resource training set is then augmented using ChatGPT-4o under various data augmentation multipliers. These multipliers determine how many times the original number of responses is increased. For example, if the multiplier is 3 and we start with 5 responses, the augmented dataset will contain 15 responses. We evaluate the fine-tuned model’s performance on the test set consisting of 64 labeled responses. The calculation of M-IoU, IoU, and F2 score was each repeated in 5 experiments, meaning a total of 15 experiments were repeated for each set size, with a different random seed used for each run.